### v2

# Many-Body Dissipative Particle Dynamics Simulations of Lipid Bilayers with the MDPD-MARTINI Force-Field

Abstract
Many-body dissipative particle dynamics
(MDPD) offers a significant speed-up in the simulation of various systems, including soft matter, in comparison with molecular dynamics (MD) simulations based on Lennard-Jones interactions, which is crucial for describing phenomena characterized by large time and length scales. Moreover, it has recently been shown that the MARTINI force-field coarse-graining approach is applicable in MDPD, thus rendering feasible the simulation of complex molecules as in MD-MARTINI for ever larger systems and longer physical times. Here, simulations of various lipid membranes were performed by using the MDPD-MARTINI coarse-grained (CG) force-field (FF), relevant properties were calculated, and comparison with standard MD-MARTINI CG simulations and experimental data was made. Thus, insights into structural properties of these bilayer systems and further evidence regarding the transferability of the MDPD-MARTINI CG models are provided. In this regard, this is a natural next step in the development of the general-purpose MDPD-MARTINI-FF, which generally provides significant speed-ups in both computational and physical simulated times, in comparison with standard CG MD simulations.
Keywords:
Many-Body Dissipative Particle Dynamics MDPD-MARTINI-FF Lipid Bilayers Water
journal: Eur. Phys. J. Plus
1 Introduction
Molecular dynamics (MD) Rapaport_2004 is a common method used for the simulation of a whole range of materials in various areas, such as soft matter, fluid dynamics, and solid-state physics. A key element for reliable MD simulations is the ability of describing the interactions between particles (force field) in such a way that various experimental properties of interest of a system are well-reproduced by the simulation. To this end, the non-bonded interactions between particles play a key role in reproducing these properties and much of the research for creating the various force-fields (FFs) is mainly focused on adequately describing those non-bonded interactions for a given cutoff at a specific temperature (the development of FFs valid over a wide range of temperatures is an active field of research), with further refinements coming from the topology FF parameters, which often can also be crucial. When interactions are considered between all atoms of the system, that is in the case of all-atom FFs, simulations are generally computationally expensive, albeit often necessary, for example, when a finer time- and length-scale resolution is required, what has led to the development of coarse-grained (CG) FFs Noid2013 ; Kmiecik2016 ; Jin2022 ; Noid2023 . In this case, interactions between groups of atoms are modelled instead of individual atoms, thus the number of required force (interactions) calculations during the integration of the equations of motion in MD simulations is significantly reduced, which, in turn, results in faster simulations and significant savings in energy resources consumed to run the simulations and cool down the hardware. Moreover, CG models offer the possibility of using a larger time step in the simulation and faster dynamics, which provides an additional speed-up of the simulations.
MARTINI is one of the most popular CG FFs for MD simulation, since it is a general-purpose FF that can be applied in the simulation of a wide spectrum of components Marrink2004 ; Marrink2007 ; Marrink2013 ; Alessandri2021 , such as proteins Monticelli2008 ; Periole2009 ; deJong2013 ; Poma2017 , polymers Lee2009 , carbohydrates Lopez2009 , glycolipids Lopez2013 , glycans Chakraborty2021 , DNA Uusitalo2015 , RNA Uusitalo2017 , water Yesylevskyy2010 and various solvents Vainikka2021 . In addition, various extensions include simulations for specific pH Grunewald2020 and more recently chemical reactions (reactive MARTINI) Sami2023 . While the MARTINI 3.0 CG FF Souza2021 has recently been released, the development of any FF is a continuous effort that aims at improving the FF parameters through refinement based on a wider range of methodologies and validation with experimental data on an ever growing number of possible systems. In the case of CG FFs, validation can often take place by comparison to results obtained by well-established all-atom FFs or CG models, which have already been parametrized according to experimental data.
Regarding the MD-MARTINI FF, its popularity is mainly due to the fact that it allows for the simulation of a wide spectrum of different systems by only using a finite number of CG bead types and a top-down approach for non-bonded interactions. Here, the concept of a ’LEGO’ approach of combining these beads to build more complex molecules while accounting for chemical specificity is used. In addition, MARTINI is based on the Lennard-Jones potential with the smallest possible cutoff including attractions, which, in principle, renders this model computationally efficient, while various features implemented in GROMACS software (e.g. virtual atoms) Berendsen1995 , namely the native platform for MARTINI, allow for further speed-up of the simulations. Currently, the MD MARTINI method enjoys a large community, which applies and further tests the models in a range of diverse systems, thus practically participating in the validation effort of this general-purpose FF.
Despite the wide use of the MD-MARTINI model, MD remains a computationally expensive method even for CG models. However, it has recently been shown that the MARTINI approach can be applied in the case of many-body dissipative particle dynamics (MDPD) Carnevale2024_MDPD_MARTINI , a method that offers a computational speed-up of about an order of magnitude in comparison with the MD method due to the short-range interactions, as well as an additional speed-up from the use of soft interactions, while at the same time it accounts for both repulsion and attraction between particles Pagonabarraga2001 ; warren2003 ; zhao2017 ; zhao2021-review ; zhao2021 ; Han2021 . In particular, it has been shown that the MDPD simulations are 4–7 times faster than the MD simulations for systems of DPPC lipids and water Carnevale2024_MDPD_MARTINI . Other comparisons in the literature have shown that MDPD models are four times faster than MD MARTINI models Zhu2021 , while in simpler systems (e.g. only water) MDPD models perform even faster Ghoufi2011 . Moreover, there is no limitation in combining MDPD and MD models in the same simulation. However, such a scenario is in general undesirable, since it will naturally lead to a higher computational cost for the simulation. Here, we build on previous work that dealt with the study of DPPC bilayers and further apply the MDPD-MARTINI-FF in the investigation of POPC and DOPC lipid bilayers and compare the obtained results with those of standard MD-MARTINI simulations. Such a comparison between the MD- and MDPD-MARTINI FFs is realized on the basis of a number of related properties for bilayer systems. Moreover, a new interaction level is provided to account for the C3 MARTINI bead-type in DOPC and POPC lipids. The MDPD-MARTINI models provide equivalent results to the MD MARTINI models, while they offer at the same time a significant speed-up in the simulations, namely several minutes of MDPD simulations instead of several hours of MD simulations, while interactions are transferable for these systems as in the case of the MD-MARTINI-FF.
In the following, we present the MDPD- and MD-MARTINI models for the various lipid systems and details on the simulation methods. Then, we present the analysis and discussion of various properties. In the last section, we draw our conclusions and discuss next steps in the development of the general-purpose MDPD-MARTINI FF.
2 Simulation models and methods
Figure 1: MD and MDPD models of DPPC, DOPC, and POPC lipids. The type of beads are indicated by the different color. In particular, $Q_{0}$ represents the choline headgroup, $Q_{a}$ the phosphate headgroup, two $N_{a}$ beads correspond to the glycerol ester moiety, while the two alkane chains consist of $C_{1}$ and $C_{3}$ (for POPC and DOPC) beads. The POPC has a double bond at carbon atom positions 9 and 10, while the DOPC has a double bond at each chain, as illustrated. The interactions between the beads are given in Table 2.
2.1 Molecular Dynamics
The MD model follows that of previous work for DPPC bilayers Carnevale2024_MDPD_MARTINI , which has been a benchmark system in the case of the MD-MARTINI-FF lipids_tutorial and therefore further details will be here omitted. Initial files for LAMMPS software Plimpton1995 ; Thompson2022 for the DPPC, DOPC, and POPC bilayers based on the MD-MARTINI model can be found as moltemplate examples lipids_tutorial_moltemplate . The DPPC model consists of different bead types, that is $Q_{0}$ for the choline headgroup, $Q_{a}$ for the phosphate headgroup, two $N_{a}$ beads for the glycerol ester moiety, and two chains, each comprised of four $C_{1}$ beads, to represent the alkane tails. For DOPC and POPC lipids, the model differs from that of DPPC lipids. This distinction arises due to the presence of double bonds in the alkane chains, with one double bond in POPC and two in DOPC. These double bonds require a different angle in the chain structure and an additional bead type, C3, to accurately represent their molecular geometry. The lipid models are illustrated in Figure 1, with the names of the bead types being the same for both the MD and MDPD MARTINI models following the MARTINI FF naming convention. Hence, this implies that both models use the same mapping of CG beads to chemical groups of atoms. Finally, water is represented by the P4 bead type.
Figure 2: Typical snapshots of DPPC, DOPC, and POPC bilayers of 512 lipids in water. Upper panel corresponds to results obtained by MD-MARTINI FF simulations, while the lower panel refers to simulated structures obtained via the MDPD-MARTINI force-field.
Typical simulations for forming the lipid bilayers follow the same protocol as in previous work Carnevale2024_MDPD_MARTINI . However, here, the aim is not to show the computational speed-up of the method during the bilayer formation, since this is well-established for MDPD simulations and becomes even apparent during the calculation of the system properties after the system has reached equilibrium Carnevale2024_MDPD_MARTINI , but rather to assess the quality performance of the MDPD-MARTINI model in the simulation of DOPC and POPC bilayers with an additional C3 bead-type and investigate key properties of these bilayers. Typically, the simulated systems consist of 512 lipids, which are placed in a simulation box of dimensions $L_{x}=L_{y}=130$ and $L_{z}=100$, maintaining a 16:1 proportion of water beads.
Preparing the bilayers involves minimizing the system’s energy, conducting a simulation in the NPT ensemble employing the Nosé–Hoover thermostat at 300 K and 1 bar to obtain the equilibrium box volume. Then, simulations are carried out at an elevated temperature (typically around 450 K) to facilitate the removal of any undesired structures that may have formed, and the final bilayer formation. Further refinement of the system can take place by realizing NPT simulations employing a zero surface tension condition while gradually lowering the temperature to 300 K. Equilibrium properties are calculated by NPT runs at this temperature with typical snapshots for each system presented in Figure 2.
2.2 Many-body dissipative particle dynamics
The MDPD method has been applied for fluids with different properties Espanol1995 ; warren2003 ; zhao2017 ; zhao2021-review ; zhao2021 ; Han2021 ; Vanya2018 ; Carnevale2023 ; Ng2025 , as well as a range of soft matter multi-phase and multi-component systems, such as systems with surfactant molecules Hendrikse2023 ; Carnevale2024 . The most common use of MDPD simulation is usually carried out by solving the Langevin equation of motion for each particle (Equation 1) as is also done in the case of MD simulations Theodorakis2011 . However, in the case of MDPD, forces are directly defined, instead of being provided by a potential form. The positions and velocities of each particle evolve according to Equation 1
$m\frac{d\bm{v}_{i}}{dt}=\sum_{j\neq i}\bm{F}_{ij}^{C}+\bm{F}_{ij}^{R}+\bm{F}_{ij}^{D},$ (1)
which considers the interactions of a particular particle $i$ with all surrounding particles within a cutoff distance, $\bm{F}^{C}$, and the presence of dissipative, $\bm{F}^{D}$, and random forces, $\bm{F}^{R}$, to act as a thermostat. The mass, $m$, is the same for all particles and set to unity.
The total conservative force on particle $i$, namely, $\bm{F}^{C}$, which is the main difference between dissipative particle dynamics (DPD) and MDPD Espanol2017 ; Yoshimoto2013 ; Li2014 ; Lavagnini2021 ; Trofimov2005 is mathematically expressed by the following relation
$\bm{F}^{C}_{ij}=A\omega^{C}(r_{ij})\bm{e}_{ij}+B\left(\bar{\rho_{i}}+\bar{\rho_{j}}\right)\omega^{d}(r_{ij})\bm{e}_{ij}$ (2)
and includes attractive ($A<0$) and repulsive ($B>0$) interactions. $r_{ij}$ is the distance between particles, $\bm{e}_{ij}$ the unit vector connecting particles $i$ and $j$, while $\omega^{C}(r_{ij})$ and $\omega^{d}(r_{ij})$ are linear weight functions given as
$\omega^{C}(r_{ij})=\begin{cases}&1-\frac{r_{ij}}{r_{c}},\ \ r_{ij}\leq r_{c}\\ &0,\ \ r_{ij}>r_{c}.\end{cases}$ (3)
$r_{c}$ is a cutoff distance for the interactions, which is usually set to unity. Moreover, $\omega^{d}(r_{ij})$ has the same form, but its cutoff distance is $r_{d}=0.75$. The repulsive term in the conservative force depends on the local densities, which are expressed as follows:
$\bar{\rho_{i}}=\sum_{0<r_{ij}\leq r_{d}}\frac{15}{2\pi r_{d}^{3}}\left(1-\frac{r_{ij}}{r_{d}}\right)^{2}.$ (4)
The random and dissipative forces are
$\bm{F}^{D}_{ij}=-\gamma\omega^{D}(r_{ij})(\bm{e}_{ij}\cdot\bm{v}_{ij})\bm{e}_{ij},$ (5)
$\bm{F}^{R}_{ij}=\xi\omega^{R}(r_{ij})\theta_{ij}\Delta t^{-1/2}\bm{e}_{ij},$ (6)
respectively, with $\gamma$ being the dissipative strength, $\xi$ the strength of the random force, $\bm{v}_{ij}$ the relative velocity between particles, and $\theta_{ij}$ a random variable from a Gaussian distribution with unit variance. The fluctuation–dissipation theorem requires that $\gamma$ and $\xi$ be related by the following relation
$\gamma=\frac{\xi^{2}}{2k_{B}T}.$ (7)
The temperature of the system in the simulations was $T=1$ (MDPD units) while the weight functions for the dissipative and random forces are
$\omega^{D}(r_{ij})=\left[\omega^{R}(r_{ij})\right]^{2}=\begin{cases}\left(1-\frac{r_{ij}}{r_{c}}\right)^{2},&r_{ij}\leq r_{c}\\ 0,&r_{ij}>r_{c}.\end{cases}$ (8)
The equations of motion (Equation 1) are integrated by a modified velocity-Verlet algorithm in LAMMPS Plimpton1995 ; Thompson2022 with a time step $\Delta t=0.01$. The non-bonded interactions between beads are tuned by varying $A_{ij}$, while repulsive interactions will remain constant as $B=35$ warren2013 . In the MDPD model, harmonic bond and angle interactions are employed, which are mathematically expressed as
$E_{bond}=k\left(r_{ij}-r_{0}\right)^{2}$ (9)
and
$E_{angle}=k_{A}\left(\theta_{ijk}-\theta_{0}\right)^{2},$ (10)
respectively. The parameters for the bond interactions are $k=150$ and $r_{0}=0.5$, while for the angles $k_{A}=5$, while $\theta_{0}$ depends on the specific molecule with values indicated in Table 1.
Angle $\theta_{0}$
$Q_{0}-Q_{a}-N_{a}$ I
$Q_{a}-N_{a}-N_{a}$ II
$Q_{a}-N_{a}-C_{1}$ I
$N_{a}-N_{a}-C_{1}$ II
$N_{a}-C_{1}-C_{3}$ I
$N_{a}-C_{1}-C_{1}$ I
$C_{1}-C_{3}-C_{1}$ II
$C_{3}-C_{1}-C_{1}$ I
$C_{1}-C_{1}-C_{1}$ I
Table 1: $\theta_{0}$ values for equilibrium angles for triads of consecutive CG beads: $180^{\circ}$ (I); $120^{\circ}$ (II).
The MDPD model for the DPPC, POPC, and DOPC is illustrated in Figure 1. Moreover, the interactions between the different beads are presented in Table 2 Carnevale2024_MDPD_MARTINI . Also, an additional bead-type C3 with respect to the DPPC lipid model has been considered as in the MD-MARTINI model in the case of the DOPC and POPC lipids. Finally, forming the bilayers by using the MDPD-MARTINI model follows the same simulation protocol as in the MD-MARTINI simulations with temperature increase from 300 K to 450 K. Typical snapshots for each bilayer case obtained from MDPD simulations are presented in Figure 2 alongside the corresponding MD-MARTINI snapshots.
P4 Q0 Qa Na $C_{1}$ $C_{3}$
P4 I I I II VI V
Q0 I IV IV II VI V
Qa I IV IV II VI V
Na II II II III VI V
$C_{1}$ VI VI VI VI VI VI
$C_{3}$ V V V V VI VI
Table 2: Interaction matrix for parametrized MDPD beads organized in six interaction levels (I-VI) with corresponding attractive parameters (A): $-50$ (I); $-43$ (II); $-34$ (III); $-30$ (IV); $-28$ (V); $-26$ (VI).
3 Results and Discussion
Figure 3: Comparison of bilayer thickness in MD (a, b) and MDPD (c, d) models. Results for POPC (a, c) and DOPC (b, d) bilayers consisting of 512 lipid molecules are shown. The values of bilayer thickness are reported in Table 3.
Table 3: MDPD and MD measurements of area per lipid and bilayer thickness for the different lipid molecules considered. Both methods produce membranes of comparable thickness, while the current MDPD force-field underestimates the APL.
MDPD MD
DPPC APL $46.2$2 $51.6$2
Thickness $44.2$ $44.0$
POPC APL $54.3$2 $61.6$2
Thickness $40.1$ $39.5$
DOPC APL $59.1$2 $64.7$2
Thickness $37.8$ $38.4$
The MD and MDPD MARTINI systems typically consisted of 512 lipids, while in certain cases, for example, for determining the spectrum intensities of the undulation modes on the bilayer, larger systems have been used (8192 lipids in total). Moreover, different sizes were tested in our previous work on DPPC lipids and all cases were able to form a stable bilayer Carnevale2024_MDPD_MARTINI . Firstly, the area per lipid (APL) was examined. This can be determined by the area a lipid covers in such bilayers, in other words, this is the area of the bilayer divided by the average number of lipids in a single monolayer. In our study, the area of a plane aligned tangentially to the bilayer surface in the simulation box was divided by half the total number of lipid molecules in order to calculate the APL. This property depends on various factors, such as the lipid type and temperature. In this case, the influence of the lipid type was examined and a comparison between the MD and MDPD models was performed. The exact values of APL for the lipid bilayers considered are reported in Table 3. In particular, we can observe that MDPD exhibits lower APL values than those from MD simulations. Moreover, both simulation methods present an increase in APL when going from DPPC–POPC–DOPC. This is expected partly due to the different angle involving the C3-type beads (Table 1).
Another fundamental property that has been estimated is the thickness bilayer for the different lipids for both the MDPD- and MD-MARTINI models. This parameter is essential for describing structural aspects of the bilayers and was determined by identifying the peaks in the density distribution profile of the choline head-group bead (Q0) as shown in Figure 3. It is found that both the MD and MDPD models are in agreement with each other and moreover with experimental data drabik2020 for all lipid bilayer cases. The exact values are reported in Table 3. In particular, it has been found that the thickness decreases as the APL increases for the different systems. The structural characteristics expressed through the density profiles of each bilayer for the various bead types are similar for both the MD and the MDPD models, as shown in Figure 3.
Figure 4: Spectrum intensitites of the undulation modes on the bilayer. Long wavelengths obey the $q^{-4}$ behavior and transition around an 1 nm wavelength to a constant value in both the MD and the MDPD simulations is observed for POPC (a) and DOPC (bilayers). Data are obtained for a system of 8192 lipids.
Figure 5: Second-rank order-parameter, $P_{2}$, calculated for every bond along the lipid molecule in the case of POPC and DOPC lipids. Each bilayer consisted of 8192 lipids.
Lipid bilayers possess elastic properties, which can be expressed by the bending modulus, $k_{c}$. In particular, this reflects the intensity of undulations within the lipid membrane and can be obtained by the fluctuation spectrum tarazona2013 . By replicating the bilayers in the bilayer plane and running short simulations to reach equilibrium structures, the undulation spectrum was determined. Specifically, undulations can be described through a surface function, $u(x,y)$, which represents the average position of the two monolayers based on the positions of the phosphate head beads. Then, the undulation spectrum, $\hat{u}$, can be obtained by performing a Fourier transform on the surface function, that is:
$\hat{u}(\mathbf{q})=\frac{1}{N}\int u(\mathbf{x})e^{-i\mathbf{q}\cdot\mathbf{x}}\mathbf{dx}.$ (11)
Moreover, the undulation intensities in the Fourier space are expected to follow the relation tarazona2013
$\langle\hat{u}^{2}\rangle=\frac{k_{B}T}{A}\left(\frac{1}{k_{c}q^{4}}+\frac{1}{\gamma q^{2}}\right),$ (12)
where $A=L_{x}L_{y}$ is the lipid-membrane area, while $q=|\mathbf{q}|=\sqrt{q_{x}^{2}+q_{y}^{2}}$ is the wavevector length in the radial direction and $\gamma$ the surface tension. Our data, shown in Figure 4, for the DOPC and POPC bilayers exhibit the $q^{-4}$ scaling for the long wavelength modes, with a very good match between the MD and the MDPD results. Moreover, in comparison with previous results for DPPC bilayers Carnevale2024_MDPD_MARTINI , here, the $q^{-2}$ dependence is not present due to the application of the zero surface-tension condition. The $q^{-4}$ dashed line was plotted by considering the experimentally obtained bending modulus $k_{c}=5^{-20}$ (rawicz2000, ).
To further compare the MD- and MDPD-MARTINI results, the second-rank order parameter $P_{2}$ was computed Chremos2016 , given by the following relation:
$P_{2}=\frac{1}{2}\left(3\langle\cos^{2}\theta\rangle-1\right)$ (13)
Here, $\theta$ is the angle between each bond and the bilayer normal direction. Thus, it describes the alignment of the lipid molecules. In particular, values closer to $P_{2}=1$ indicate a higher degree of alignment with the normal direction, whereas values around $P_{2}=0$ rather point to a random orientation. Finally, values towards $P_{2}=-0.5$ indicate anti-alignment. The obtained results indicate that the MDPD-MARTINI model can reproduce the alignment characteristics of the lipid molecules as in the MD simulations.
4 Conclusions
The MDPD-MARTINI-FF was employed for investigating structural properties of lipid bilayers with results being compared to simulations based on the widely used MD-MARTINI CG FF. The two models are in very good agreement in a range of different properties discussed in this study, thus providing further support in favor of using MDPD models, since simulations are much faster in terms of computational and physical times Carnevale2024_MDPD_MARTINI . This particular feature opens the door for reaching larger systems and longer times scales in the simulation of complex systems, covering the full-range of chemical groups currently available in the MD-MARTINI models. This will also have the potential of accelerating the acquisition of physical-based data, which can be used in machine learning or other artificial-intelligence methods. In various applications, the presence of soft interactions in MDPD models may pose certain limitations when hard-core interactions might be necessary. While this remains to be seen, various solutions can be envisioned in such a scenario beyond the possible use of hybrid models based on MDPD and LJ interactions. In addition, further work with MDPD models is expected to provide more information regarding any limitations of the MDPD models for specific systems, as MDPD models gain in popularity.
In this study, the FF parameters of the C3 MARTINI bead-type has been provided to the MDPD-MARTINI interaction matrix Carnevale2024_MDPD_MARTINI . In view of the transferability of the potential based on the hydrophobic/hydrophilic characteristics of chemical groups represented by the CG beads, which is an important concept in the MARTINI coarse-graining approach and demonstrated here in the case of DPPC, POPC, and DOPC lipids, a range of different systems can be simulated by the provided set of interactions. Thus, by following the MARTINI mapping of CG beads to chemical units additional systems can be investigated for further validations. Finally, methods, such as the GōMARTINI for proteinsPoma2017 , will be integrated into the general-purpose MDPD-MARTINI-FF as things progress.
Acknowledgments
This research has been supported by the National Science Centre, Poland, under grant
No. 2019/34/E/ST3/00232. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016608.
Data Availability Statement
Data sets generated during the current study are available from the corresponding author on reasonable request.